OpenStudy (anonymous):

Prove: 1+sinx/1-sinx=csc^2x+2cscx+1

OpenStudy (anonymous):

Please Help! It's really important

OpenStudy (anonymous):

Welcome to Openstudy

OpenStudy (anonymous):

Just remember that \(\mathbb {cscx=\frac{1}{sinx}}\) and use \((a-b)(a+b)=a^2-b^2\)

OpenStudy (anonymous):

is the proof solvable because iv'e tried everything?

OpenStudy (phi):

if you mean (1+sinx)/(1-sinx)=csc^2x+2cscx+1 this is not an identity. It is true for specific values of x i.e. you can solve for the x's that make it true.

OpenStudy (anonymous):

how could i solve for x

OpenStudy (phi):

Up above you said Prove: Is this supposed to be an identity that they want you to prove ? (though it's not an identity, this may be a question with a "typo" and they meant to give an identity)

OpenStudy (phi):

Or do they want you to "solve for x" ?

OpenStudy (anonymous):

It's supposed to be a trig identity that he wants us to prove

OpenStudy (anonymous):

I have been trying for hours to prove it

OpenStudy (anonymous):

I tried and found out that this is not an identity. @phi is right

OpenStudy (anonymous):

If the right side is \(sec^2x+2secx+1\), then it will be an identity. Is that right, @phi?

OpenStudy (phi):

no, I don't think its sec^2 + 2sec + 1 (but if you can show it works, I'm all ears) I'm trying to figure out what they might have meant, but so far, no luck.

OpenStudy (anonymous):

I will try

OpenStudy (phi):

First, is it really \[ \frac{1+\sin x}{1-\sin x}= \csc^2 x+2\csc x+1\]?

OpenStudy (anonymous):

It says, PROVE: you can manipulate the left side to make it 'look' like the right side. Do not manipulate the right side.

OpenStudy (anonymous):

yes that is what the question is

OpenStudy (phi):

what do you get if you let x=30 degrees for both sides ?

OpenStudy (anonymous):

Yeah, @phi is right. Neither way is an identity

OpenStudy (phi):

sin(30)= 0.5 (1+0.5)/(1-0.5) = 3 csc(x)= 1/sin(x) so csc(30) = 2 csc^2(30)= 4 the right side is csc^2(30) + 2 csc(30) + 1 = 4 +2 +1 = 7 3 does not equal 7 that proves it's not an identity

OpenStudy (anonymous):

Okay! Thank you so much for your help!

OpenStudy (phi):

**the right side is csc^2(30) + 2 csc(30) + 1 = 4 +2*2 +1 = 9 3 does not equal 9 that proves it's not an identity

OpenStudy (anonymous):

It's just weird that he would ask that question!

OpenStudy (phi):

so it might be a "mistake" or he wanted you to show it's not true.